# Infrared spectra of methane-containing ice mixtures for JWST data analysis

### v1

@@ 3.2 B335 methane survey
To test the usability of the obtained spectra we fitted the 7.7 m methane absorption band of the Class 0 protostar B335. JWST Mid-IR Instrument (MIRI) Medium Resolution Spectrograph (MRS) observations toward B335 are provided within the Investigating Protostellar Accretion proposal (PID 1802, P.I. T. Megeath; Megeath et al., 2021). Level 3 observation data were taken from the MAST database (MAST DOI: 10.17909/yr06-tf19), we further scaled the CH2-SHORT channel data to match the CH1-LONG signal level. MIRI MRS observations included three levels of JWST calibration pipeline with version 1.19.1 and Calibration References Data System (CRDS) context jwst_1413.pmap (CRDS_VER = ’12.1.11’) describing in Greenfield and Miller (2016); Bushouse et al. (2024); van Gelder et al. (2024). Unfortunately, the overlap region coincides with the OCN- absorption band and we excluded this specie from the further analysis.
@@
B335 is viewed edge-on and has two prominent outflows (e.g. Bjerkeli et al., 2019), the west of which displaying strong continuum emission around 7.7 m. Based on the intensity map at 8 m (Fig. 2) we picked multiple apertures that cover bright pixels. This approach maximizes the signal-to-noise ratio, enhancing the analysis of the finer spectral details. To date, this is the first study that explores the spatial distribution on MIRI-MRS JWST data for a molecule with relatively low abundance relative to H2O. We note that a spatial distribution study was carried out recently for solid H2O, CO and CO2 on the NIRCam data (Smith et al., 2025). The spectra were extracted from the cylindrical apertures: seven 1.53′′ apertures were used to assess the spatial distribution of methane and a 4.59′′ aperture was used to ensure consistency between fits in smaller apertures. The center positions for the chosen apertures are listed in Table 2.
@@
All spectra were processed in the same way. After baseline subtraction a linear combination of methane laboratory mixtures was fitted to the observational spectrum. The linear scale coefficients were obtained by minimizing , given as:
@@
where and are observational and modeled optical depths, respectively, and is the point-wise standard deviation of the observational spectrum. The standard deviation was obtained from the residuals between the original optical depth spectrum and a smoothed curve obtained with locally weighted scatterplot smoothing (LOWESS, Cleveland, 1979) with a 24-point window. Column densities of H2O and CO2 were estimated based on 6 m and 15.2 m bands, respectively.
@@
The fitted spectra are shown in the Fig. 3. The - observational spectra correspond to the - apertures from Fig. 2. Fitting results show that about 70% and 30% of methane budget in B335 are best described by methane in H2O and methane in CO2 environments, respectively (see Table 2). A minor part of methane is found in CH3OH environment, however, the detection at this level may be sensitive to e.g. baseline selection. Spectrum is excluded from the analysis due to low signal-to-noise ratio. The 1.53′′ apertures display variations in ice composition with CH4 in CO2 to CH4 in H2O relative abundances decreasing towards the protostar. There is also an overall increase in column density towards the core, likely explained by an increase of density towards the center.
@@
The presence of CH4 in CO2 matrix provides some insights on ice formation at early stages of the protostar. An interesting correlation is observed between CH4 and CO2. The Spearman’s and p-values (Spearman, 1904) for CO2 and CH4 column densities (, =0.007) and abundances (, =0.003) indicate strong, statistically significant, positive correlation between these species. The correlations between CH4 and CO2 are known from as early as Spitzer methane survey (Öberg et al., 2008). Modeling presented in Garrod and Pauly (2011) shows that under certain conditions methane and CO2 can form concurrently, occupying the innermost layers of ice mantles. Ferrero et al. (2024) explored the reactivity of 3C-OH2 complex with various species, revealing barrierless reactions with atomic O and H. These reactions lead to production of CH2 and CO, which are the direct precursors to CH4 and CO2, respectively. While parallel formation of methane and CO2 is not largely explored, these results can be interpreted as evidence towards the concurrence for these processes.
@@ 4 Astrochemical implications and Conclusions
In this paper we present spectra of pure methane and binary mixtures with methane for 6.7 K and 10 K deposition temperatures.
@@
The following mixtures are presented in this paper: CH4:H2O = 1:10, CH4:CO2 = 1:5, CH4:CO2 = 1:15, CH4:CH3OH = 1:3 and CH4:NH3 = 1:3. All spectra of methane in astrochemically relevant mixture ratios are presented for the first time at 6.7 K. The new band strength values and spectra obtained at temperature below 9 K reveal the following astrochemical implications:
@@
1. Currently, the band strength of pure methane at 10 K is used majorly for column density estimations in observational data (e.g. McClure et al., 2023; Rocha et al., 2024, 2025). However, solid interstellar methane is found in mixtures with other icy species rather than in pure form. We also note that interstellar ices differ from laboratory analogues both in how they are produced and how they are processed. Thus, while the band strengths in Table 1 provide a better basis for interpreting observations, they are still approximate. Nevertheless, our new data invites a discussion on recent literature estimates. Since the methane band strengths in mixtures are 20 % higher on average than the band strength of pure methane at 10 K, the column densities of interstellar methane may be overestimated by this value in e.g. McClure et al. (2023). Interestingly, this partially explains the underestimation of solid methane abundance by astrochemical models toward background stars NIR38 and J110621 reported in Jiménez-Serra et al. (2025).
@@
2. We present the the first publicly available digital spectral library of methane absorbance spectra in non-H2O environment, providing the possibility to identify non-H2O methane environments on newly obtained JWST spectra. Solid methane is typically considered to reside in polar water-rich environment due its early formation along with water ice (see e.g. Hasegawa et al., 1992; Clément et al., 2023; Borshcheva et al., 2025). This is further supported by successful fits of methane features in JWST spectra of interstellar ices with CH4:H2O mixtures (e.g. Chen et al., 2024; Rocha et al., 2024, 2025). However, there are possible evidences for its presence in mixtures with other molecules. For example, works by Boogert et al. (1996) and Nakibov et al. (2025) show that observed shapes of 7.7 m methane feature can be partially interpreted with non-H2O methane-containing mixtures. Additionally, icy methane mixed with apolar species such as N2 and CO is found on icy bodies in the outer Solar system (e.g., Emtiaz et al., 2022). The obtained spectra of ices of astrochemical ratios are published on Zenodo (DOI: 10.5281/zenodo.15350566). Comparison between spectra obtained at 6.7 K and 10 K reveal temperature-induced changes in shape and peak position, although we find it unlikely that these differences are prominent enough to be discerned within the observational data. However, it is known that central regions of prestellar cores are well-shielded from external heating radiation. Consequently, as shown both observationally (Pagani et al., 2004, 2007; Crapsi et al., 2007; Pagani et al., 2015) and theoretically (Keto and Field, 2005; Keto and Caselli, 2010, and references therein), dust temperature there may reach 6 K. According to a recent modeling by Borshcheva et al. (2025), approximately 80% of solid methane in the prestellar core L1544 resides on grains with temperature below 9 K. Thus, to accurately interpret future observations of ices on a line of sight toward L1544, laboratory spectra and band strengths of mixtures obtained both at 6.7 K and 10 K might be needed. The presented spectra can be directly compared to observational data, because the deformation mode of methane is unlikely to be affected by grain size, depending mostly on temperature and molecular environment (Boogert et al., 1997).
@@
3. We present the first spatial survey performed on JWST MIRI-MRS data, exploring methane in B335. The results show that about 30% of solid methane is found in CO2 environment, rather than in H2O matrix. The consistency between findings in full and smaller apertures proves the usability of presented spectra for JWST data analysis. The observed ice composition gradient further highlights that JWST provides enough sensitivity for such studies, prompting further research.
5 Data availability
@@
The laboratory infrared absorbance spectra in transmission regime in the 2.5–15.9 m (4000–630 cm-1) region for CH4 in different environments at 6.7 K and 10 K can be found on Zenodo at the following link: https://zenodo.org/records/15350566.
@@ Acknowledgements.
The authors thank Prof. Dr. Paola Caselli for the discussion on the dust temperature in prestellar cores, and Ms. Katerina Borshcheva for providing the information on the distribution of methane ice in the model of L1544. We thank the anonymous reviewer for their comments that helped us to improve the manuscript. This research work is funded by the Russian Science Foundation via 23-12-00315 agreement.References
@@
There are three phases of solid methane under vacuum conditions. The crystalline phase II is stable below 20.4 K and has the fcc lattice with eight sublattices: six of which have site symmetry D2d and two — site symmetry Oh (Kobashi et al. 1977). It displays a double peak structure of both (3009 cm-1 and 3021 cm-1) and (1301 cm-1 and 1298 cm-1) features. This phase reversibly transitions into phase I above the first transition temperature of 20.4 K and has fcc lattice with a single freely rotating molecule. The and features are positioned at 3010 cm-1 and 1299 cm-1, respectively. Phase II* is observed if methane ice is obtained by vapor deposition below 10 K. It displays a degree of orientational disorder between phase I and phase II (Emtiaz et al. 2020). In this phase the absorption bands of methane are positioned near 3010 cm-1 and 1297 cm-1 (Gerakines and Hudson 2015b). Phase II* cannot be obtained by cooling methane ice from the triple point or from phase II ice. On warm-up phase II* ice undergoes irreversible transition to phase II. The second transition temperature depends on experimental conditions due to metastability and is observed below 10 K (Colwell et al. 1962; Krupskii et al. 1973; Emtiaz et al. 2020).
@@
Despite the extensive research carried out on methane solid phases in infrared at low temperatures in recent decades (Kobashi et al. 1977; Hudgins et al. 1993; Boogert et al. 1997; Mulas et al. 1998; Gerakines and Hudson 2015b; Hudson et al. 2015; Emtiaz et al. 2020; Gerakines and Hudson 2020), there is still an ambiguity between published studies. In several works (Hudgins et al. 1993; Gerakines et al. 1996; Mulas et al. 1998; Alsindi et al. 2003; de Barros et al. 2011) methane is claimed to be in phase II, but the presented spectral features differ significantly from the theoretically derived shape (Kobashi et al. 1977). Gerakines and Hudson (2015b) addressed this discrepancy by providing spectra of ‘amorphous methane’, suggesting that the literature spectra in the works above display partial amorphous-to-crystalline transition, which was caused by the high deposition rates and experimental constraints of the setups used.
@@
Gerakines and Hudson (2015b) report that the majority of their spectra were obtained in the 8–12 K range, but the spectra are referred to as obtained ‘at 10 K’ for simplicity throughout the text. Thus, it is not clear at what temperature exactly were published experiments carried out, a point that also applies to Hudson et al. (2015). This ambiguity is significant, because the 8–12 K range includes published methane second transition temperatures (Colwell et al. 1962; Krupskii et al. 1973). We attempted to obtain the spectrum at their Fig. 1(f) at precisely 10 K, but failed even using a lower deposition rate than stated in their paper. Therefore, prior to obtaining the set of methane-bearing mixtures at 10 K we decided to investigate the transition between phase II* and phase II. Under the deposition rate and duration of experiments described in the main text methane retained phase II* structure when deposited at 6.7 K throughout the whole deposition time. In contrast, for deposition at 10 K, the double peak structure characteristic of phase II started to form once ice reached certain column density, see Figure 4. At a deposition rate of 5.9 cm-2 s-1 the transition started after 35 minutes, corresponding to a column density of 1.21016 cm-2. We also conducted depositions at rates twice as low and twice as high: 2.95 cm-2 s-1 and 1.18 cm-2 s-1. In all cases, on further deposition the double peak evolved through intermediate shapes, similar to those shown by Gerakines and Hudson (2015b), before fully converting to the phase II structure. The change in peak structure occurs at a critical value of 1.21016 cm-2. This equals approximately to 6 monolayers of pure methane ice on each side of the Ge substrate, given that single monolayer column density is 11015 cm-2.
@@
From our experiments we conclude that phase II* is stable at 6.7 K. At 10 K phase II* is only observed in very thin methane films. In our experiments on Ge substrate, the maximum ice column density at which phase II* is still observed at 10 K is 1.21016 cm-2, independent of the deposition rate. Further ice growth at 10 K leads to the transition of phase II* to phase II through intermediate spectral shapes. While these intermediate spectra are similar to those mentioned in Gerakines and Hudson (2015b) we refrain from a more direct comparison. We expect that similar results can be obtained under minor differences in experimental conditions of other setups, which is expected given the environment-sensitive nature of metastable phase transition. A definitive answer would come from X-ray diffractometry study as performed in Mizuno et al. (2016, 2022), which showed that other molecular ices that are ‘conventionally amorphous’ for astrochemistry can exhibit positionally ordered but orientationally disordered structure (CO or N2), while other (CO2, CS2) are indeed amorphous.
@@
We also note an instrumental effect that concerns both temperature programmed desorption (TPD) and static temperature experiments with volatile species. Here, we interpret the cases transparently reported by Collings et al. (2004) and Schrauwen et al. (2025). Methane temporarily adsorbs on warm surfaces (some cryostat parts, radiation shield, etc.) during the deposition and continuously releases even when the leak valves are closed, as in Schrauwen et al. (2025). The evacuation therefore is governed by the amount of temporarily stored methane rather than by turbomolecular pump’s efficiency, resulting in a ‘tail’ in QMS signal after deposition. During TPD warm surfaces release stored material early, polluting the QMS signal, as in Collings et al. (2004). Released gas can get readsorbed by the cold substrate. An example of that can be found in the cyclic experiment in Gerakines and Hudson (2015b) (see their Fig. 1, right column). We note that the integrated intensity increased after the first heating-holding-cooling cycle, indicating an increase in column density. Second cycle did not cause an increase as stored methane depleted on first heating. This increase did not affect the results of their paper.

### v2

@@ 3.2 B335 methane survey
To test the usability of the obtained spectra, we fitted the 7.7 m methane absorption band of the Class 0 protostar B335. JWST Mid-IR Instrument (MIRI) Medium Resolution Spectrograph (MRS) observations toward B335 are provided within the Investigating Protostellar Accretion proposal (PID 1802, P.I. T. Megeath; Megeath et al., 2021). Level 3 observation data were taken from the Mikulski Archive for Space Telescopes (MAST) database (DOI: 10.17909/yr06-tf19), and we further scaled the CH2-SHORT channel data to match the CH1-LONG signal level. MIRI MRS observations included three levels of the JWST calibration pipeline with version 1.19.1 and Calibration References Data System (CRDS) context jwst_1413.pmap (CRDS_VER = ’12.1.11’), which is described in Greenfield and Miller (2016), Bushouse et al. (2024), and van Gelder et al. (2024). Unfortunately, the overlap region coincides with the OCN- absorption band, and we excluded this species from further analysis.
@@
The object B335 is viewed edge-on and has two prominent outflows (e.g., Bjerkeli et al., 2019), the western of which displays strong continuum emission around 7.7 m. Based on the intensity map at 8 m (Fig. 2), we selected multiple apertures that covered bright pixels. This approach maximizes the signal-to-noise ratio, which facilitates our analysis of the finer spectral details. This is the first study that explores the spatial distribution on MIRI-MRS JWST data for a molecule with a low abundance relative to H2O. We note that a spatial distribution study was carried out recently for solid H2O, CO, and CO2 on NIRCam data (Smith et al., 2025). The spectra were extracted from the cylindrical apertures: seven 1.53′′ apertures were used to assess the spatial distribution of methane, and a 4.59′′ aperture was used to ensure consistency between fits in smaller apertures. The center positions for the chosen apertures are listed in Table 2.
@@
All spectra were processed in the same way. After baseline subtraction, a linear combination of methane laboratory mixtures was fitted to the observational spectrum. The linear scale coefficients were obtained by minimizing , given as
@@
where and are observational and modeled optical depths, respectively, and is the point-wise standard deviation of the observational spectrum. The standard deviation was obtained from the residuals between the original optical depth spectrum and a smoothed curve obtained with locally weighted scatterplot smoothing (Cleveland, 1979) with a 24-point window. The column densities of H2O and CO2 were estimated based on the 6 m and 15.2 m bands, respectively.
@@
The fitted spectra are shown in Fig. 3. The - observational spectra correspond to the - apertures from Fig. 2. The fitting results show that about 70% and 30% of the methane budget in B335 are best described by methane in H2O and methane in CO2 environments, respectively (see Table 2). A minor fraction of the methane is found in the CH3OH environment, but the detection at this level may be sensitive to baseline selection, for example. Spectrum was excluded from the analysis because the signal-to-noise ratio was low. The 1.53′′ apertures display variations in ice composition, with CH4 in CO2 to CH4 in H2O relative abundances decreasing toward the protostar. The column density toward the core also increases overall, which is likely explained by an increase in density toward the center.
@@
The presence of CH4 in the CO2 matrix provides some insights into ice formation in early stages of the protostar. An interesting correlation is observed between CH4 and CO2. The Spearman and p values (Spearman, 1904) for the CO2 and CH4 column densities (, =0.007) and abundances (, =0.003) indicate a strong statistically significant positive correlation between these species. The correlations between CH4 and CO2 are known from the Spitzer methane survey (Öberg et al., 2008). The modeling presented in Garrod and Pauly (2011) shows that under certain conditions, methane and CO2 can form concurrently, occupying the innermost layers of ice mantles. Ferrero et al. (2024) explored the reactivity of the 3C-OH2 complex with various species, revealing barrierless reactions with atomic O and H. These reactions lead to the production of CH2 and CO, which are the direct precursors of CH4 and CO2, respectively. While the parallel formation of methane and CO2 has not been largely explored, these results can be interpreted as evidence that these processes occur.
@@ 4 Astrochemical implications and conclusions
We have presented the spectra of pure methane and binary mixtures with methane for 6.7 K and 10 K deposition temperatures: CH4:H2O = 1:10, CH4:CO2 = 1:5, CH4:CO2 = 1:15, CH4:CH3OH = 1:3, and CH4:NH3 = 1:3. All spectra of methane in astrochemically relevant mixture ratios are presented for the first time at 6.7 K. The new band strength values and spectra obtained at a temperature below 9 K revealed the astrochemical implications listed below.
@@
1. The band strength of pure methane at 10 K is currently mainly used for column density estimations in observational data (e.g., McClure et al., 2023; Rocha et al., 2024, 2025). However, solid interstellar methane is found in mixtures with other icy species rather than in pure form. We also note that interstellar ices differ from laboratory analogs both in how they are produced and in how they are processed. Thus, while the band strengths in Table 1 provide a better basis for interpreting observations, they are still approximate. Nevertheless, our new data invite a discussion on recent literature estimates. Since the methane band strengths in mixtures are 20 % higher on average than the band strength of pure methane at 10 K, the column densities of interstellar methane may be overestimated by this value by McClure et al. (2023), for instance. Interestingly, this partially explains the underestimation of the solid methane abundance by astrochemical models toward the background stars NIR38 and J110621 reported in Jiménez-Serra et al. (2025).
@@
2. We have presented the the first publicly available digital spectral library of methane absorbance spectra in a non-H2O environment, providing the possibility to identify non-H2O methane environments in newly obtained JWST spectra. Solid methane is typically considered to reside in a polar water-rich environment because it forms early along with water ice (see, e.g., Hasegawa et al., 1992; Clément et al., 2023; Borshcheva et al., 2025). This is further supported by successful fits of methane features in JWST spectra of interstellar ices with CH4:H2O mixtures (e.g., Chen et al., 2024; Rocha et al., 2024, 2025). However, there is possible evidence that it is present in mixtures with other molecules. For example, Boogert et al. (1996) and Nakibov et al. (2025) show that the observed shapes of the 7.7 m methane feature can be partially interpreted with non-H2O methane-containing mixtures. Additionally, icy methane mixed with apolar species such as N2 and CO is found on icy bodies in the outer Solar System (e.g., Emtiaz et al., 2022). The obtained spectra of ices of astrochemical ratios are published on Zenodo (DOI: 10.5281/zenodo.15350566). Our comparison of spectra obtained at 6.7 K and 10 K revealed temperature-induced changes in shape and peak position, although we find it unlikely that these differences are prominent enough to be discerned within the observational data. However, it is known that the central regions of pre-stellar cores are well shielded from external heating radiation. Consequently, as shown both observationally (Pagani et al., 2004, 2007; Crapsi et al., 2007; Pagani et al., 2015) and theoretically (Keto and Field, 2005; Keto and Caselli, 2010, and references therein), the dust temperature there can reach 6 K. According to recent modeling by Borshcheva et al. (2025), approximately 80% of solid methane in the pre-stellar core L1544 resides on grains with temperatures below 9 K. Thus, to accurately interpret future observations of ices on a line of sight toward L1544, laboratory spectra and band strengths of mixtures obtained at both 6.7 K and 10 K might be needed. The presented spectra can be directly compared to observational data because the deformation mode of methane is unlikely to be affected by grain size; it mostly depends on temperature and the molecular environment (Boogert et al., 1997).
@@
3. We have presented the first spatial survey performed on JWST MIRI-MRS data and explored methane in B335. The results show that about 30% of solid methane is found in a CO2 environment and not in the H2O matrix. The consistency between the findings from full and smaller apertures proves that the spectra we have presented can be used to analyze JWST data. The observed ice composition gradient further demonstrates that JWST provides enough sensitivity for such studies, which should motivate further research.
Data availability
@@
The laboratory IR absorbance spectra in the transmission regime in the 2.5–15.9 m (4000–630 cm-1) region for CH4 in different environments at 6.7 K and 10 K can be found on Zenodo via the following link: https://zenodo.org/records/15350566.
@@ Acknowledgements.
The authors thank Prof. Dr. Paola Caselli for the discussion on the dust temperature in prestellar cores, and Ms. Katerina Borshcheva for providing the information on the distribution of methane ice in the model of L1544. We thank the anonymous reviewer for their comments that helped us to improve the manuscript and A&A Language Editor Astrid Peter for corrections in the English text. This research work is funded by the Russian Science Foundation via 23-12-00315 agreement.References
@@ Appendix A Pure CH4
There are three phases of solid methane under vacuum conditions. The crystalline phase II is stable below 20.4 K and has the fcc lattice with eight sub-lattices: six of which have site symmetry D2d and two of which have site symmetry Oh (Kobashi et al. 1977). It displays a double peak structure of both (3009 cm-1 and 3021 cm-1) and (1301 cm-1 and 1298 cm-1) features. This phase reversibly transitions into phase I above the first transition temperature of 20.4 K and has fcc lattice with a single freely rotating molecule. The and features are positioned at 3010 cm-1 and 1299 cm-1, respectively. Phase II* is observed if methane ice is obtained by vapor deposition below 10 K. It displays a degree of orientational disorder between phase I and phase II (Emtiaz et al. 2020). In this phase the absorption bands of methane are positioned near 3010 cm-1 and 1297 cm-1 (Gerakines and Hudson 2015b). Phase II* cannot be obtained by cooling methane ice from the triple point or from phase II ice. On warming phase II* ice undergoes irreversible transition to phase II. The second transition temperature depends on experimental conditions due to metastability and is observed below 10 K (Colwell et al. 1962; Krupskii et al. 1973; Emtiaz et al. 2020).
@@
Despite the extensive research carried out on methane solid phases in the IR at low temperatures in recent decades (Kobashi et al. 1977; Hudgins et al. 1993; Boogert et al. 1997; Mulas et al. 1998; Gerakines and Hudson 2015b; Hudson et al. 2015; Emtiaz et al. 2020; Gerakines and Hudson 2020), there is still an ambiguity between published studies. In several works (Hudgins et al. 1993; Gerakines et al. 1996; Mulas et al. 1998; Alsindi et al. 2003; de Barros et al. 2011) methane is claimed to be in phase II, but the presented spectral features differ significantly from the theoretically derived shape (Kobashi et al. 1977). Gerakines and Hudson (2015b) addressed this discrepancy by providing spectra of ‘amorphous methane’, suggesting that the literature spectra in the works above display partial amorphous-to-crystalline transition, which was caused by the high deposition rates and experimental constraints of the setups used.
@@
Gerakines and Hudson (2015b) report that the majority of their spectra were obtained in the 8–12 K range, but the spectra are referred to as obtained ‘at 10 K’ for simplicity throughout the text. Thus, it is not clear at what temperature exactly the published experiments were carried out, a point that also applies to Hudson et al. (2015). This ambiguity is significant, because the 8–12 K range includes published methane second transition temperatures (Colwell et al. 1962; Krupskii et al. 1973). We attempted to obtain the spectrum in their Fig. 1f at precisely 10 K, but failed even using a lower deposition rate than stated in their paper. Therefore, prior to obtaining the set of methane-bearing mixtures at 10 K we decided to investigate the transition between phase II* and phase II. Under the deposition rate and duration of experiments described in the main text methane retained phase II* structure when deposited at 6.7 K throughout the whole deposition time. In contrast, for deposition at 10 K, the double peak structure characteristic of phase II started to form once the ice reached certain column density; see Fig. 4. At a deposition rate of 5.9 cm-2 s-1 the transition started after 35 minutes, corresponding to a column density of 1.21016 cm-2. We also conducted depositions at rates twice as low and twice as high: 2.95 cm-2 s-1 and 1.18 cm-2 s-1. In all cases, on further deposition the double peak evolved through intermediate shapes, similar to those shown by Gerakines and Hudson (2015b), before fully converting to the phase II structure. The change in peak structure occurs at a critical value of 1.21016 cm-2. This equals approximately to 6 monolayers of pure methane ice on each side of the Ge substrate, given that a single monolayer column density is 11015 cm-2.
@@
From our experiments we conclude that phase II* is stable at 6.7 K. At 10 K phase II* is only observed in very thin methane films. In our experiments on Ge substrate, the maximum ice column density at which phase II* is still observed at 10 K is 1.21016 cm-2, independent of the deposition rate. Further ice growth at 10 K leads to the transition of phase II* to phase II through intermediate spectral shapes. While these intermediate spectra are similar to those mentioned in Gerakines and Hudson (2015b) we refrain from a more direct comparison. We expect that similar results can be obtained under minor differences in experimental conditions of other setups, which is expected given the environment-sensitive nature of metastable phase transition. A definitive answer would come from X-ray diffractometry study as performed in Mizuno et al. (2016, 2022), which showed that other molecular ices that are ‘conventionally amorphous’ for astrochemistry can exhibit positionally ordered but orientationally disordered structure (CO or N2), while others (CO2, CS2) are indeed amorphous.
@@
We also note an instrumental effect that concerns both temperature programmed desorption (TPD) and static temperature experiments with volatile species. Here, we interpret the cases transparently reported by Collings et al. (2004) and Schrauwen et al. (2025). Methane temporarily adsorbs on warm surfaces (some cryostat parts, radiation shield, etc.) during the deposition and continuously releases even when the leak valves are closed, as in Schrauwen et al. (2025). The evacuation therefore is governed by the amount of temporarily stored methane rather than by the turbomolecular pump’s efficiency, resulting in a ‘tail’ in the QMS signal after deposition. During TPD warm surfaces release stored material early, polluting the QMS signal, as in Collings et al. (2004). Released gas can get re-adsorbed by the cold substrate. An example of that can be found in the cyclic experiment in Gerakines and Hudson (2015b, see their Fig. 1, right column). We note that the integrated intensity increased after the first heating-holding-cooling cycle, indicating an increase in column density. The second cycle did not cause an increase because the stored methane was depleted on the first heating. This increase did not affect the results of their paper.